OpenStudy (anonymous):

How to turn 0.1857857857... (a recurring decimal) to a fraction?

OpenStudy (faiqraees):

x=0.1857857857 10x = 1.857857857 1000x = 1857.857857 1000x-10x = 1857.857... - 1.857... 990x=1856 x = 1856/990 Simplify the fraction and you have your answer

OpenStudy (anonymous):

I think I got it! Thanks!!!